OpenStudy (anonymous):

use integration by parts to find integral of (ln(4x))^2

OpenStudy (anonymous):

\[\int\limits_{}^{}(\ln(4x))^2\] dx

OpenStudy (anonymous):

I'm not sure what to make u and what to make dv...

OpenStudy (blurbendy):

you are familiar with the formula for using integration by parts, correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[uv-\int\limits_{}^{}v (du)\]

OpenStudy (blurbendy):

\[\int\limits_{}^{}fdg = fg - \int\limits_{}^{}gdf\] i would start with this f = ln^2|4x| df = (2ln|4x|)dx dg = dx g = x Can you proceed from here?

OpenStudy (blurbendy):

same formula, i just used the notation im familiar with

OpenStudy (anonymous):

okay, Yeah I think so. Thanks for the help!

OpenStudy (blurbendy):

wait df should be 2ln|4x|) / (x) dx

OpenStudy (blurbendy):

np